# Longest cycles in vertex-transitive and highly connected graphs

### v1

@@ Abstract.
We present progress on two old conjectures about longest cycles in graphs. The first conjecture, due to Thomassen from 1978, states that apart from a finite number of exceptions, all connected vertex-transitive graphs contain a Hamiltonian cycle. The second conjecture, due to Smith from 1984, states that for $r\geq 2$ in every $r$-connected graph any two longest cycles intersect in at least $r$ vertices. In this paper, we prove a new lemma about the intersection of longest cycles in a graph which can be used to improve the best known bounds towards both of the aforementioned conjectures: First, we show that every connected vertex-transitive graph on $n\geq 3$ vertices contains a cycle of length at least $\Omega(n^{13/21})$, improving on $\Omega(n^{3/5})$ from [De Vos, arXiv:2302:04255, 2023]. Second, we show that in every $r$-connected graph with $r\geq 2$, any two longest cycles meet in at least $\Omega(r^{5/8})$ vertices, improving on $\Omega(r^{3/5})$ from [Chen, Faudree and Gould, J. Combin. Theory, Ser.  B, 1998]. Our proof combines combinatorial arguments, computer-search and linear programming.
@@ 1. Introduction
One of the oldest and most fundamental research areas in graph theory is concerned with longest cycles [^1] Throughout this paper, a longest cycle in a graph will be defined as a cycle attaining the maximum possible length among all cycles in the graph. in graphs. In particular, sufficient conditions for the Hamiltonicity of a graph, that is, the existence of a cycle of length $n$ in a graph $G$ on $n$ vertices, have been widely studied, going back to the classical result of Dirac [17] stating that every graph with minimum degree at least $n/2$ is Hamiltonian for all $n\geq 3$. The minimum degree condition is tight in Dirac’s theorem, so to obtain similar results for sparser graphs one needs to add additional constraints, such as connectivity conditions. For $2$-connected regular graphs, the degree condition was improved to $n/3$ by Jackson [31] and later to the tight $n/3-1$ by Hilbig [30], answering a question of Szekeres (see [31]). For $3$-connected regular graphs, asymptotically confirming a conjecture [^2] Their conjecture was made for general $t$-connected $d$-regular graphs, $t\geq 3$ and said that $d\geq n/t+1$ is sufficient to guarantee Hamiltonicity. However, this was disproved for $t\geq 4$ by by Jung [33] and independently by Jackson, Li and Zhu [32]. independently made by Bollobás [6] and Häggkvist [29], Kühn, Lo, Osthus and Staden [35] showed that minimum degree $n/4+o(n)$ is sufficient to guarantee the graph being Hamiltonian. More generally, in the same paper they also considered some sufficient conditions on the length of the longest cycle of a graph. An old conjecture of Bondy [7], confirmed by Wei [45], says that for all $r\geq 3$, every sufficiently large $2$-connected regular graph on $n$ vertices with degree at least $n/r$ has a cycle of length at least $2n/(r-1)$. Kühn, Lo, Osthus and Staden [35] considered a generalization of this conjecture for $t$-connected graphs and showed that degree at least $n/r+o(n)$ guarantees a cycle of length at least $\min\{tn/(r-1),(n-o(n))\}$.
@@
In contrast to the dense case, sufficient conditions for Hamiltonicity that also address sparse graphs, with, say, a constant average degree, seem to be rather scarce. The positive results that apply to sparse graphs can be roughly clustered into three (mostly orthogonal) research directions: conditions related to pseudo-randomness such as expanders (see e.g. the recent breakthrough [18]); toughness conditions [14]; and symmetry conditions such as vertex-transitivity. Note that for dense vertex-transitive graphs, that is, of regularity $d=\Theta(n)$, Hamiltonicity is already established. Indeed, every connected vertex-transitive graph of degree $d$ is $32(d+1)/$ 3-connected [24], thus the results from [35] discussed above imply the existence of a cycle of length $(1-o(1))n$.
@@
In fact, even stronger results hold: Christofides, Hladkỳ and Máthé [13] showed that all connected vertex transitive graphs of regularity of order $\Theta(n)$ contain a Hamiltonian cycle. Our first main result concerns longest cycles in general connected vertex transitive graphs.
@@
There are two main open problems in the study of Hamiltonicity on vertex transitive graphs that are commonly attributed to Lovász and Thomassen, respectively. The first is a conjecture posed by Lovász [38] in 1969 and asserts that every connected vertex-transitive graph admits a Hamiltonian path. The second is a conjecture by Thomassen from 1978 (see e.g. [5], Conjecture 1 and the references therein) and states that apart from finitely many exceptional graphs, all connected vertex-transitive graphs even contain a Hamiltonian cycle. In fact, to this date only four connected vertex transitive graphs on at least $3$ vertices are known which do not have a Hamiltonian cycle (one of which is the infamous Petersen graph). Despite half a century of study, during which the conjectures of Lovász and Thomassen received much attention by the research community, they remain wide open today. In particular, for 44 years the best known general lower bound on the maximum length of a path or cycle in a connected vertex-transitive graph on $n\geq 3$ vertices remained $\Omega(\sqrt{n})$ due to Babai [4], and only recently this bound was improved to $\Omega(n^{3/5})$ by DeVos [16]. There is some further recent evidence of this conjecture being true though: indeed results of Draganić, Mongtomery, Correia, Pokrovskiy and Sudakov [18], together with an earlier results on expansion properties of random Cayley graphs obtained by Alon and Roichman [1], imply that random Cayley graphs with generator set of size at least $\Omega(\log{n})$ are Hamiltonian. For more related work, we refer to the surveys [2, 36] on the partial progress towards Lovász’ and Thomassen’s conjectures. As our first main result, we obtain the following improved lower bound for the length of longest cycles in vertex-transitive graphs.
@@
In this section we explain how to deduce both Theorems 1.1 and 1.3 from Lemma 1.4. Let us start with Theorem 1.3.
@@
Let $G$ be an $r$-connected graph, and let $C_{1},C_{2}$ be two longest cycles in $G$. The cycles $C_{1}$ and $C_{2}$ must overlap in at least one vertex, since two disjoint cycles with two vertex-disjoint paths between them always gives rise to a cycle of length strictly larger than the average of the two cycles lengths. This implies that $k:=|V(C_{1})\cap V(C_{2})|>0$. So by Lemma 1.4, a smallest vertex-separator $S$ between $V(C_{1})\setminus V(C_{2})$ and $V(C_{2})\setminus V(C_{1})$ in $G$ must have size at most $|S|\leq O(k^{8/5})$. On the other hand, since $G$ is $r$-connected, we must have $|S|\geq r$, implying $k\geq\Omega(r^{5/8})$, as desired. ∎
@@
In the following, let us define a longest cycle transversal as a set $S\subseteq V(G)$ which intersects every longest cycle of $G$. For $k\in\mathbb{N}$, let $s(k)$ denote the smallest integer $s$ such that every 2-connected graph $G$ containing two longest cycles $C_{1}$ and $C_{2}$ such that $|V(C_{1})\cap V(C_{2})|=k$ has a longest cycle transversal of size at most $s$. Note that if $G$ is vertex transitive and connected of degree $d$, then it is not only two connected but $\frac{2}{3}(d+1)$ connected [24, Theorem 3.4.2]. DeVos [16, Lemma 3] showed that $s(k)\leq k^{2}+k$, and the remainder of the proof in [16] directly shows that the length of the longest cycle in a connected vertex-transitive graph on at least $n\geq 3$ vertices is at least
@@
We next show how to obtain the improved upper bound $s(k)=O(k^{8/5})$ from Lemma 1.4. Given two longest cycles $C_{1}$ and $C_{2}$ in a $2$-connected graph $G$ with $|V(C_{1})\cap V(C_{2})|=k$, by Lemma 1.4 there exists a vertex-separator $S$ of size at most $O(k^{8/5})$ between $V(C_{1})\setminus V(C_{2})$ and $V(C_{2})\setminus V(C_{1})$ in $G$. We now claim that $S\cup(V(C_{1})\cap V(C_{2}))$ (which is of size $\leq|S|+k=O(k^{8/5})$) intersects all longest cycles in $G$. Indeed, for any longest cycle $C$ in $G$, since $G$ is $2$-connected, we must have $V(C)\cap V(C_{i})\neq\emptyset$ for $i=1,2$. Hence, either $C$ contains a vertex in $V(C_{1})\cap V(C_{2})$ or a segment of $C$ forms a path from $V(C_{1})\setminus V(C_{2})$ to $V(C_{2})\setminus V(C_{1})$ in $G$, in which case this segment (and thus $C$) must be hit by $S$. This shows that indeed, $s(k)\leq O(k^{8/5})$.
@@ 3. Reduction to computer search
While reading the definitions in this section, the reader may wish to keep the following definition of a bipartite graph in mind as motivation. Given a graph $G$ with two longest cycles $C_{1}$ and $C_{2}$ and a collection of vertex-disjoint paths $\mathcal{P}$ from $C_{1}-V(C_{2})$ to $C_{2}-V(C_{1})$, we may define a bipartite graph as follows. The vertices are the connected components of $C_{1}-V(C_{2})$ and $C_{2}-V(C_{1})$, and there is an edge between two such components if there is a vertex-disjoint path $P$ between them with $P\in\mathcal{P}$.
@@
Given an ordered bipartite graph $\mathcal{H}=(H,\sigma_{X},\sigma_{Y})$, an $\mathcal{H}$-configuration is a tuple $\tau=(\tau_{h})_{h\in V(H)}$, where $\tau_{h}$ is a total order on the neighbours of $h$ for each $h\in V(H)$. To this configuration, we create an associated configuration graph, denoted by $G_{\tau}$, which we define as follows. First we create a vertex $v_{x,y}$ for each $x\in X$ and $y\in N_{H}(x)$ and let $C_{1}^{\tau}:=\{v_{x,y}:x\in X,~{}y\in N_{H}(x)\}$. We turn this into a cycle using the orders $(\sigma_{X},(\tau_{x})_{x\in X})$ “lexicographically”. That is, we consider the total order $\leq_{C_{1}^{\tau}}$ obtained by
$v_{x_{1},y_{1}}\leq_{C_{1}^{\tau}}v_{x_{2},y_{2}}$ if $\quad\text{if $x_{1}\leq_{\sigma_{X}}x_{2}$},$
$v_{x_{1},y_{1}}\leq_{C_{1}^{\tau}}v_{x_{1},y_{2}}$ if $\quad\text{if $y_{1}\leq_{\tau_{x_{1}}}y_{2}$}.$
We place edges to turn $C_{1}^{\tau}$ into a cycle [^3] If $C_{1}^{\tau}$ has only one vertex, this yields a loop in the graph, and if $C_{1}^{\tau}$ has two vertices, this creates parallel edges., adding an edge between the smallest and largest vertex in this order and otherwise adding edges between vertices that are consecutive in the total order. Although $C_{1}^{\tau}$ is undirected, we also fix a cyclic order on it, which arises naturally by following the order described above. We define $C_{2}^{\tau}$ analogously by switching the roles of $X$ and $Y$. Finally, we define $G_{\tau}$ as the union of $C_{1}^{\tau}$ and $C_{2}^{\tau}$ together with the edges $\{\{v_{x,y},v_{y,x}\}:\{x,y\}\in E(H)\}\}$. We will call these “edges corresponding to $H$” the cross edges. We call $C_{1}^{\tau}$ and $C_{2}^{\tau}$ the main cycles of $G_{\tau}$.
@@
The next lemma will be useful for checking for the existence of long good cycles independent of the weight function.
@@
Let $\mathcal{H}=(H,\sigma_{X},\sigma_{Y})$ be an ordered bipartite graph and let $\tau$ be an $\mathcal{H}$-configuration. If there are two good cycles $C$ and $C^{\prime}$ in $G_{\tau}$ which together contain all of the edges of its main cycles, as well as at least one cross edge, then for any weight function $w$, $(G_{\tau},w)$ contains a long good cycle.
@@
We are only interested in the feasibility, so we may set the objective function to 1. This linear program is not feasible (there are no solutions) if and only if for every weight function $w$, $(G_{\tau},w)$ contains a long good cycle. We record this fact below.
@@
Let $\mathcal{H}=(H,\sigma_{X},\sigma_{Y})$ be an ordered bipartite graph and let $\tau$ be an $\mathcal{H}$-configuration. Then LPH,τ is infeasible if and only if for all weight functions $w$, $(G_{\tau},w)$ contains a long good cycle.
@@
  • Let $i\in\{1,2\}$ and suppose that $v,w\in V(C_{i}^{\tau^{\prime}})\subseteq V(C_{i}^{\tau})$ satisfy $vw\in E(C_{i}^{\tau^{\prime}})$. If $P$ is the unique path [^4] When $C_{1}^{\tau^{\prime}}$ has only two vertices $x$ and $x^{\prime}$, we distinguish “the edge from $x$ to $x^{\prime}$” and “the edge from $x^{\prime}$ to $x$” to obtain a unique path here. When $C_{1}^{\tau^{\prime}}$ contains only one vertex $x$, this path refers to the loop on $x$. between these vertices in $C_{i}^{\tau}$ that contains no other vertices of $C_{i}^{\tau^{\prime}}$, then $w(P)=w^{\prime}(vw)$.
We are now ready to show that it is sufficient to find long good cycles in subconfigurations.
@@
Let $\mathcal{H}^{\prime}$ and $\mathcal{H}$ denote ordered subgraphs, $\tau^{\prime}$ an $\mathcal{H}^{\prime}$-configuration and $\tau$ an $\mathcal{H}$-configuration. Suppose that $\tau^{\prime}$ is a subconfiguration of $\tau$ and $w$ is a weight function on the configuration graph $G_{\tau}$ which extends a weight function $w^{\prime}$ on the configuration graph $G_{\tau^{\prime}}$. If $(G_{\tau^{\prime}},w^{\prime})$ contains a long good cycle, then $(G_{\tau},w)$ also contains a long good cycle.
@@
Let $\mathcal{H}=(H,\sigma_{X},\sigma_{Y})$ be an ordered bipartite graph such that either $Q_{3}^{+}$ or $K_{3,3}$ is a subgraph of $H$, and let $\tau$ be an $\mathcal{H}$-configuration. For any weight function $w$ on $G_{\tau}$, $(G_{\tau},w)$ contains a long good cycle.
@@
We then run a computer search to check that for every configuration $\tau$ on these three ordered bipartite graphs, every weighted configuration graph resulting from $\tau$ contains a long good cycle. As there are many configurations even on these three ordered bipartite graphs, we proceed by adding edges iteratively. Indeed, it is a direct consequence of Lemma 3.3 that is it sufficient to check the condition for at least one subconfiguration of every possible configuration.
@@
    At step $i\in\{1,\dots,|E(H)|\}$, we construct $S_{i}$ as follows. For each $(H_{i-1},\sigma_{X},\sigma_{Y})$-configuration $\tau^{\prime}\in S_{i-1}$, consider every $(H_{i},\sigma_{X},\sigma_{Y})$-configuration $\tau$ such that $\tau^{\prime}$ is a sub-configuration of $\tau$. Let $R_{i}$ be the set of these configurations. Add $\tau$ to $S_{i}$ if the criterion of Lemma 3.1 cannot guarantee that every weighted configuration graph on $\tau$ would contain a long good cycle.
@@
    Finally, we create a set $S$ as follows. For every $H$-configuration $\tau$ in $S_{|E(H)|}$, we add $\tau$ to $S$ if the more powerful linear program criterion from 3.2 cannot guarantee that every weighted configuration graph on $\tau$ would contain a long good cycle.

### v2

@@ Abstract.
We present progress on three old conjectures about longest paths and cycles in graphs. The first pair of conjectures, due to Lovász from 1969 and Thomassen from 1978, respectively, states that all connected vertex-transitive graphs contain a Hamiltonian path, and that all sufficiently large such graphs even contain a Hamiltonian cycle. The third conjecture, due to Smith from 1984, states that for $r\geq 2$ in every $r$-connected graph any two longest cycles intersect in at least $r$ vertices. In this paper, we prove a new lemma about the intersection of longest cycles in a graph which can be used to improve the best known bounds towards all the aforementioned conjectures: First, we show that every connected vertex-transitive graph on $n\geq 3$ vertices contains a cycle (and hence path) of length at least $\Omega(n^{13/21})$, improving on $\Omega(n^{3/5})$ from [DeVos, arXiv:2302:04255, 2023]. Second, we show that in every $r$-connected graph with $r\geq 2$, any two longest cycles meet in at least $\Omega(r^{5/8})$ vertices, improving on $\Omega(r^{3/5})$ from [Chen, Faudree and Gould, J. Combin. Theory, Ser.  B, 1998]. Our proof combines combinatorial arguments, computer-search and linear programming.
@@ 1. Introduction
One of the oldest and most fundamental research areas in graph theory is concerned with longest cycles [^1] Throughout this paper, a longest cycle in a graph will be defined as a cycle attaining the maximum possible length among all cycles in the graph. in graphs. In particular, sufficient conditions for the Hamiltonicity of a graph, that is, the existence of a cycle of length $n$ in a graph on $n$ vertices, have been widely studied, going back to the classical result of Dirac [19] stating that every graph with minimum degree at least $n/2$ is Hamiltonian for all $n\geq 3$. The minimum degree condition is tight in Dirac’s theorem, so to obtain similar results for sparser graphs one needs to add additional constraints, such as connectivity conditions. For $2$-connected regular graphs, the degree condition was improved to $n/3$ by Jackson [33] and later to the tight n/31n/3-1 by Hilbig [32], answering a question of Szekeres (see [33]). For $3$-connected regular graphs, asymptotically confirming a conjecture [^2] Their conjecture was made for general $t$-connected $d$-regular graphs, $t\geq 3$ and said that $d\geq n/t+1$ is sufficient to guarantee Hamiltonicity. However, this was disproved for $t\geq 4$ by by Jung [35] and independently by Jackson, Li and Zhu [34]. independently made by Bollobás [8] and Häggkvist [31], Kühn, Lo, Osthus and Staden [37] showed that minimum degree $n/4+o(n)$ is sufficient to guarantee the graph being Hamiltonian. More generally, in the same paper they also considered some sufficient conditions on the length of the longest cycle of a graph. An old conjecture of Bondy [9], confirmed by Wei [47], says that for all $r\geq 3$, every sufficiently large $2$-connected regular graph on $n$ vertices with degree at least $n/r$ has a cycle of length at least 2n/(r1)2n/(r-1). Kühn, Lo, Osthus and Staden [37] considered a generalization of this conjecture for $t$-connected graphs and showed that degree at least $n/r+o(n)$ guarantees a cycle of length at least $\min\{tn/(r-1),(n-o(n))\}$.
@@
In contrast to the dense case, sufficient conditions for Hamiltonicity that also address sparse graphs, with, say, a constant average degree, seem to be rather scarce. The positive results that apply to sparse graphs can be roughly clustered into three (mostly orthogonal) research directions: conditions related to pseudo-randomness such as expanders (see e.g. the recent breakthrough [20]); toughness conditions [7, 16]; and symmetry conditions such as vertex-transitivity. Note that for dense vertex-transitive graphs, that is, of degree $d=\Theta(n)$, a lower bound of (1o(1))n(1-o(1))n on the maximum cycle length follows from the aforementioned results: Indeed, every connected vertex-transitive graph of degree $d$ is $2(d+1)/3$-connected [26], and thus the results from [37] discussed above imply the existence of a cycle of length (1o(1))n(1-o(1))n.
@@
In fact, even stronger results hold: Christofides, Hladkỳ and Máthé [15] showed that all connected vertex-transitive graphs of degree of order $\Theta(n)$ contain a Hamiltonian cycle. Our first main result concerns longest cycles in general connected vertex-transitive graphs.
@@
There are two main open problems in the study of Hamiltonicity on vertex-transitive graphs that are commonly attributed to Lovász and Thomassen, respectively. The first is a conjecture posed by Lovász [40] in 1969 and asserts that every connected vertex-transitive graph admits a Hamiltonian path. The second is a conjecture by Thomassen from 1978 (see e.g. [6], Conjecture 1 and the references therein) and states that apart from finitely many exceptional graphs, all connected vertex-transitive graphs even contain a Hamiltonian cycle. In fact, to this date only four connected vertex-transitive graphs on at least $3$ vertices are known which do not have a Hamiltonian cycle (one of which is the infamous Petersen graph). It is worth mentioning that Lovász originally posed his conjecture in the opposite form, but nowadays the above mentioned positive version has become standard. Curiously, Babai [5] has proposed a conjecture contradicting Lovász conjecture, namely that there exists an absolute constant $\varepsilon>0$ such that for infinitely many $n$ there exists a vertex-transitive graph of order $n$ with no path of length larger than $(1-\varepsilon)n$. Despite half a century of study, during which the conjectures of Lovász and Thomassen received much attention by the research community, they remain wide open today. In particular, for 44 years the best known general lower bound on the maximum length of a path or cycle in a connected vertex-transitive graph on $n\geq 3$ vertices remained $\Omega(\sqrt{n})$ due to Babai [4], and only recently this bound was improved to $\Omega(n^{3/5})$ by DeVos [18]. There is some further recent evidence of this conjecture being true though: indeed results of Draganić, Montgomery, Munhá Correia, Pokrovskiy and Sudakov [20], together with earlier results on expansion properties of random Cayley graphs obtained by Alon and Roichman [1], imply that random Cayley graphs with generator set of size at least $\Omega(\log{n})$ are Hamiltonian. For more related work, we refer to the surveys [2, 38] on the partial progress towards Lovász’ and Thomassen’s conjectures. As our first main result, we obtain the following improved lower bound for the length of longest cycles in vertex-transitive graphs.
@@
In this section we explain how to deduce both Theorems 1.1 and 1.3 from Lemma 1.4. We start by recording the following folklore-observation about intersections of longest paths and cycles in graphs, which we will need repeatedly, see for example [51] for a reference.
Observation 2.1.
Let $G$ be a connected graph. Then any two longest paths in $G$ meet in some vertex. Additionally, if $G$ is $2$-connected, then any two longest cycles in $G$ meet in some vertex.
Let us now continue by first proving Theorem 1.3.
@@
Let $G$ be an $r$-connected graph, and let $C_{1},C_{2}$ be two longest cycles in $G$. Since $r\geq 2$ by assumption, Observation 2.1 implies that the cycles $C_{1}$ and $C_{2}$ must overlap in at least one vertex. This implies that $k:=|V(C_{1})\cap V(C_{2})|>0$. So by Lemma 1.4, a smallest vertex-separator $S$ between $V(C_{1})\setminus V(C_{2})$ and $V(C_{2})\setminus V(C_{1})$ in $G$ must have size at most $|S|\leq O(k^{8/5})$. On the other hand, since $G$ is $r$-connected, we must have $|S|\geq r$, implying $k\geq\Omega(r^{5/8})$, as desired. ∎
@@
In the following, let us define a longest cycle transversal as a set $S\subseteq V(G)$ which intersects every longest cycle of $G$. For $k\in\mathbb{N}$, let $s(k)$ denote the smallest integer $s$ such that every 2-connected graph $G$ containing two longest cycles $C_{1}$ and $C_{2}$ such that $|V(C_{1})\cap V(C_{2})|=k$ has a longest cycle transversal of size at most $s$. DeVos [18, Lemma 3] showed that $s(k)\leq k^{2}+k$, and the remainder of the proof in [18] directly shows that the length of the longest cycle in a connected vertex-transitive graph on at least $n\geq 3$ vertices is at least
@@
We next show how to obtain the improved upper bound $s(k)=O(k^{8/5})$ from Lemma 1.4. Given two longest cycles $C_{1}$ and $C_{2}$ in a $2$-connected graph $G$ with $|V(C_{1})\cap V(C_{2})|=k$, by Lemma 1.4 there exists a vertex-separator $S$ of size at most $O(k^{8/5})$ between $V(C_{1})\setminus V(C_{2})$ and $V(C_{2})\setminus V(C_{1})$ in $G$. We now claim that $S\cup(V(C_{1})\cap V(C_{2}))$ (which is of size $\leq|S|+k=O(k^{8/5})$) intersects all longest cycles in $G$. Indeed, for any longest cycle $C$ in $G$, since $G$ is $2$-connected and since by Observation 2.1 any two longest cycles in a $2$-connected graph must intersect each other, we must have $V(C)\cap V(C_{i})\neq\emptyset$ for $i=1,2$. Hence, either $C$ contains a vertex in $V(C_{1})\cap V(C_{2})$ or a segment of $C$ forms a path from $V(C_{1})\setminus V(C_{2})$ to $V(C_{2})\setminus V(C_{1})$ in $G$, in which case this segment (and thus $C$) must be hit by $S$. This shows that indeed, $s(k)\leq O(k^{8/5})$.
@@ 3. Reduction to computer search
While reading the definitions in this section, the reader may wish to keep the following construction of an auxiliary bipartite graph in mind as motivation. Given a graph $G$ with two longest cycles $C_{1}$ and $C_{2}$ and a collection of vertex-disjoint paths $\mathcal{P}$ from $C_{1}-V(C_{2})$ to $C_{2}-V(C_{1})$, we may define a bipartite graph as follows. The vertices are the connected components of $C_{1}-V(C_{2})$ and $C_{2}-V(C_{1})$, and there is an edge between two such components if there is a vertex-disjoint path $P$ between them with $P\in\mathcal{P}$.
@@
Given an ordered bipartite graph $\mathcal{H}=(H,\sigma_{X},\sigma_{Y})$, an $\mathcal{H}$-configuration is a tuple $\tau=(\tau_{h})_{h\in V(H)}$, where $\tau_{h}$ is a total order on the neighbours of $h$ for each $h\in V(H)$. Associated with this configuration, we create what we call a configuration graph, denoted by $G_{\tau}$, defined as follows. First we create a vertex $v_{x,y}$ for each $x\in X$ and $y\in N_{H}(x)$ and let $C_{1}^{\tau}:=\{v_{x,y}:x\in X,~y\in N_{H}(x)\}$. We turn this into a cycle using the orders $(\sigma_{X},(\tau_{x})_{x\in X})$ “lexicographically”. That is, we consider the total order $\leq_{C_{1}^{\tau}}$ obtained by
$v_{x_{1},y_{1}}\leq_{C_{1}^{\tau}}v_{x_{2},y_{2}}$ $\quad\text{if $x_{1}<_{\sigma_{X}}x_{2}$},$
$v_{x_{1},y_{1}}\leq_{C_{1}^{\tau}}v_{x_{1},y_{2}}$ $\quad\text{if $y_{1}\leq_{\tau_{x_{1}}}y_{2}$}.$
We place edges to turn $C_{1}^{\tau}$ into a cycle [^3] If $C_{1}^{\tau}$ has only one vertex, this yields a loop in the graph, and if $C_{1}^{\tau}$ has two vertices, this creates parallel edges., adding an edge between the smallest and largest vertex in this order and otherwise adding edges between vertices that are consecutive in the total order. Although $C_{1}^{\tau}$ is undirected, we also fix a cyclic order on it, which arises naturally by following the order described above. We define $C_{2}^{\tau}$ analogously by switching the roles of $X$ and $Y$. Finally, we define $G_{\tau}$ as the union of $C_{1}^{\tau}$ and $C_{2}^{\tau}$ together with the edges $\{\{v_{x,y},v_{y,x}\}:\{x,y\}\in E(H)\}\}$. We will call these “edges corresponding to HH” the cross edges. We call $C_{1}^{\tau}$ and $C_{2}^{\tau}$ the main cycles of $G_{\tau}$.
@@
The next lemma will be useful for checking for the existence of good long cycles independent of the weight function.
@@
Let $\mathcal{H}=(H,\sigma_{X},\sigma_{Y})$ be an ordered bipartite graph and let $\tau$ be an $\mathcal{H}$-configuration. If there are two good cycles $C$ and $C^{\prime}$ in $G_{\tau}$ which together contain all of the edges of its main cycles, as well as at least one cross edge, then for any weight function $w$, $(G_{\tau},w)$ contains a good long cycle.
@@
We are only interested in the feasibility, so we may set the objective function to 1. This linear program is not feasible (there are no solutions) if and only if for every weight function $w$, $(G_{\tau},w)$ contains a good long cycle. We record this fact below.
@@
Let $\mathcal{H}=(H,\sigma_{X},\sigma_{Y})$ be an ordered bipartite graph and let $\tau$ be an $\mathcal{H}$-configuration. Then LPH,τ is infeasible if and only if for all weight functions $w$, $(G_{\tau},w)$ contains a good long cycle.
Proof.
By definition of LPH,τ and by definition of a weight function, an edge-indexed vector $(w_{e})_{e\in E(G_{\tau})}$ is feasible for LPH,τ if and only if (1) the corresponding mapping $w:E(G_{\tau})\rightarrow\mathbb{R}$, defined as $w(e):=w_{e}$ for every $e$, is a weight function, and (2) we have $w(C)=\sum_{e\in E(C)}w_{e}\leq\sum_{e\in E(C_{1}^{\tau})}w_{e}=w(C_{1}^{\tau})$ for every good cycle $C$. The latter is equivalent to saying that there exist no good long cycles. Hence, it follows that LPH,τ is feasible if and only if there exists a weight function $w$ such that $(G_{\tau},w)$ contains no good long cycle. The claim of the lemma now follows by negating bot sides of this equivalence. ∎
@@
  • Let $i\in\{1,2\}$ and suppose that $v,w\in V(C_{i}^{\tau^{\prime}})\subseteq V(C_{i}^{\tau})$ satisfy $vw\in E(C_{i}^{\tau^{\prime}})$. If $P$ is the unique path [^4] When $C_{1}^{\tau^{\prime}}$ has only two vertices $x$ and $x^{\prime}$, we distinguish “the edge from $x$ to $x^{\prime}$” and “the edge from $x^{\prime}$ to xx” to obtain a unique path here. When $C_{1}^{\tau^{\prime}}$ contains only one vertex $x$, this path refers to the loop on $x$. between these vertices in $C_{i}^{\tau}$ that contains no other vertices of $C_{i}^{\tau^{\prime}}$, then $w(P)=w^{\prime}(vw)$.
We are now ready to show that it is sufficient to find good long cycles in subconfigurations.
@@
Let $\mathcal{H}^{\prime}$ and $\mathcal{H}$ denote ordered subgraphs, $\tau^{\prime}$ an $\mathcal{H}^{\prime}$-configuration and $\tau$ an $\mathcal{H}$-configuration. Suppose that $\tau^{\prime}$ is a subconfiguration of $\tau$ and $w$ is a weight function on the configuration graph $G_{\tau}$ which extends a weight function $w^{\prime}$ on the configuration graph $G_{\tau^{\prime}}$. If $(G_{\tau^{\prime}},w^{\prime})$ contains a good long cycle, then $(G_{\tau},w)$ also contains a good long cycle.
@@
Let $\mathcal{H}=(H,\sigma_{X},\sigma_{Y})$ be an ordered bipartite graph such that either $Q_{3}^{+}$ or $K_{3,3}$ is a subgraph of $H$, and let $\tau$ be an $\mathcal{H}$-configuration. For any weight function $w$ on $G_{\tau}$, $(G_{\tau},w)$ contains a good long cycle.
@@
We then run a computer search to check that for every configuration $\tau$ on these three ordered bipartite graphs, every weighted configuration graph resulting from $\tau$ contains a good long cycle. As there are many configurations even on these three ordered bipartite graphs, we proceed by adding edges iteratively. Indeed, it is a direct consequence of Lemma 3.3 that is it sufficient to check the condition for at least one subconfiguration of every possible configuration.
@@
    At step $i\in\{1,\dots,|E(H)|\}$, we construct $S_{i}$ as follows. For each $(H_{i-1},\sigma_{X},\sigma_{Y})$-configuration $\tau^{\prime}\in S_{i-1}$, consider every $(H_{i},\sigma_{X},\sigma_{Y})$-configuration $\tau$ such that $\tau^{\prime}$ is a sub-configuration of $\tau$. Let $R_{i}$ be the set of these configurations. Add $\tau$ to $S_{i}$ if the criterion of Lemma 3.1 cannot guarantee that every weighted configuration graph on $\tau$ would contain a good long cycle.
@@
    Finally, we create a set $S$ as follows. For every $H$-configuration $\tau$ in $S_{|E(H)|}$, we add $\tau$ to $S$ if the more powerful linear program criterion from 3.2 cannot guarantee that every weighted configuration graph on $\tau$ would contain a good long cycle.